OpenStudy (anonymous):

Choose the slope-intercept equation of the line that passes through the point (6, -6) and is perpendicular to y = 3x - 6. someone please help

OpenStudy (anonymous):

Perpendicular Slope= opposite reciprocal So for example, if the equation was y= 2/3 x -2, then the perpendicular slope would be -3/2 Use that, and then point slope form and plug the points and slope in.

OpenStudy (anonymous):

i dont get it

OpenStudy (anonymous):

Do you get the perpendicular slope part?

OpenStudy (anonymous):

slope of the line perpendicular to a line with slope m is -1/m

OpenStudy (anonymous):

so will it be -1/3 x - 4?

OpenStudy (anonymous):

"Perpendicular lines are a bit more complicated. If you visualize a line with positive slope (so it's an increasing line), then the perpendicular line must have negative slope (because it will be a decreasing line). So perpendicular slopes have opposite signs. The other "opposite" thing with perpendicular slopes is that their values are reciprocals; that is, you take the one slope value, and flip it upside down. Put this together with the sign change, and you get that the slope of the perpendicular line is the "negative reciprocal" of the slope of the original line — and two lines with slopes that are negative reciprocals of each other are perpendicular to each other. In numbers, if the one line's slope is m = 4/5, then the perpendicular line's slope will be m = –5/4. If the one line's slope is m = –2, then the perpendicular line's slope will be m = 1/2." So, the slope here is 3, so the opposite is -1/3. Slope intercept form is: y = mx + b. m=the slope, b= the y intercept. To find b, we plug: -6=(-1/3)(6)+b -6=-2+b -2 -2 -8=b So, y= mx+b \[y=-\frac{ 1 }{ 3 }x-8\] Hope this helped. :o) http://www.purplemath.com/modules/slope2.htm

OpenStudy (anonymous):

what are you some kinda wizard?

OpenStudy (anonymous):

What do you mean?